OpenStudy (anonymous):

PreCalculus Question: y=-4-cosπx graphing help?

OpenStudy (anonymous):

What kind of graphing calc do you have?

OpenStudy (anonymous):

If you've already graphed it, set your window to.... -5.25<x<5.25 -7.39<y<7.39

OpenStudy (cruffo):

\[y = A\cos\left(k(x-h)\right) + B\] Amplitude is A period is 2pi/k phase shift is h average is B You graph should be a cosine function, flipped on the x-axis, shifted down 4, and have a period of 2. Since there is no change in amplitude (other than the flip) the graph will range between -5 and -3 on the y-axis. Since there is not a phase shift, start with 0. Divide the period into 4 equal parts at x = 0, 0.5, 1, 1.5, 2. Then evaluate the function at these values. \[\begin{array}{c|c} x &y\\ \hline 0&-5\quad \text{ (lowest point)}\\ 0.5&-4\quad \text{(average)}\\ 1&-3\quad \text{ (highest point)}\\ 1.5&-4\quad \text{(average)}\\ 2&-5\quad \text{ (lowest point)}\\ \end{array}\]